A circular disc of mass 20kg and radius 15cm is mounted in an horizontal cylindrical axle of radius
1.5cm. Calculate the K.E of the disc after 1.2sec if s force 12N is applied tangentially to the axle.
1
Expert's answer
2021-09-20T14:42:21-0400

"KE=I\omega^2/2"


"I=mR^2/2=20\cdot0.15^2/2=0.225\ (kg\cdot m^2)"


"\omega=\epsilon t"


"M=I\epsilon \to \epsilon=M/I=12\cdot0.015/0.225=0.8\ (rad/s^2)"


"\omega=0.8\cdot 1.2=0.96\ (rad/s)"


"KE=I\omega^2/2=0.225\cdot0.96^2/2=0.104\ (J)" . Answer